) payiment OTncOime tax D) payment of interest Question 8 (1 point) Corporation X needs $1,000,000 and can raise this through debt at an annual rate of 6 percent, or preferred stock at an annual cost of 8 percent. If the corporation has a 21 percent tax rate, the after-tax cost of each is A) debt: $60,000; preferred stock: $80,000 B) debt: $47.400: preferred stock: $63,200 C) debt: $47.400: preferred stock: $80,000 D) debt: $60,000; preferred stock: $63,200 Question 9 (1 point) Saved The term red herring refers to A) the fact that most firms conducting an IPO are losing money, also known as running red ink B) a firm that is conducting an IPO without fully complying with all government 0 6
Answer #1

Correct option is C. Please see below..

Cost of debt = Interest (1-tax rate)

= 1,000,000 * 6% (1-0.21)

= 60,000 (0.79)

= 47,400

Cost of preferred stock = Interest expense

= 1,000,000 * 8%

= 80,000

Please note that there will be no tax element involved in preferred stock interest because it is shown below after tax income.